The mean days-to-flower for a population of orchids is 130 days. You notice a strikingly beautiful group of orchids in the population that you would like to cross for the next generation. You note that the mean days-to-flower in this group that you cross is 104 days. You would like to breed this particular line of orchids, but you want to know the heritability of specific traits, including days-to-flower. Luckily, you recorded the number of days-to-flower in the parents and offspring! You found the midparent-offspring regression produces a slope of 0.75. What is the mean number of days-to-flower in the offspring?

A. 67.1 days
B. 158.9 days
C. 101.1 days
D. 110.5 days

Researchers are working on a species of plant that is normally an obligate outcrosser (it does not self fertilize, but instead depends on pollen from a different individual). A mutation arises in the population that allows plants to self-fertilize, but with greatly reduced germination ability (so selection acts very strongly against this mutation). The researchers discover that this mutation reappears at a high frequency (10-4 per generation). What can be said about the relative frequency of this mutation in the population?

A. The allele for self-fertilization will prevent speciation in this population
B. The allele for self-fertilization will be maintained at medium frequency in the population.
C. The allele for self-fertilization will go to fixation because of the high mutation rate
D. The allele for self-fertilization will be eliminated from the population over time
Answer #1

As per Chegg’s regulation, I have only answered the 1st question. Please upload the questions separately to get an answer

As per Breeder's equation for heritability R = h2S

R = response to selection (the difference between mean trait in parents and mean trait in offspring)

h2 is mid-parent-offspring regression slope or heritability

S = Selection differential (the difference between mean trait in general population and mean trait in the population selected for breeding)

Let us assume that the mean number of days-to-flower in the offspring is X.

So, R = (130 days - X)

h2 = 0.75

S = (130 - 104) days = 26 days

130 days – X = 0.75*26days

or 130 days – X = 19.5 days

or X = (130 -19.5) days

or X = 110.5 days

The mean number of days-to-flower in the offspring is 110.5 days.

Kindly revert for any queries and concerns.